Circular No. 6139

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NOVA CENTAURI 1995 W. Liller, Instituto Isaac Newton, Viña del Mar, reports his discov-

ery of a probable nova with PROBLICOM on Feb. 23 and his CCD confirmation on Feb. 24, the latter yielding the position α = 13^{h}02^{m}32^{s}.07, δ = −60^{o}11′39′′.8 (equinox 2000.0; blend with GSC 8989.2001 at end figures 33^{s}.52, 53′′.7, mag 12.4). Magnitudes and colors: Jan. 27, [12 (Liller); Feb. 23.31 UT, 7.2 (Liller; Kodak TP film + orange filter); 24.04, V = 7.59 (Liller; CCD, comparison HD 113012, assumed V = 8.13, B − V = +0.11); 24.15, V = 7.70 (Liller); 25.03, m_{v} = 8.0 (D. Overbeek, Edenvale, South Africa); 25.41, V = 8.14, U − B = −0.68, B − V = +0.31, V − R = +1.01, V −I = +1.17 (A. C. Gilmore, Mt. John Observatory; comparisons Cousins F218 and F205, airmass 1.53).

J. B. Hearnshaw and L. C. Watson, Mt. John University Observatory, obtained several blue and red CCD spectra with the medium-resolution spectrograph on the 1-m telescope around Feb. 25.43 UT. These showed strong Balmer emission lines with P-Cyg profiles. The FWHM of the Hγ and Hδ lines was 5.5 nm.

SUPERNOVA 1995G IN NGC 1643 A. V. Filippenko, University of California at Berkeley, reports that D.

Schlegel obtained CCD spectra (range 335–730 nm, resolution 0.7 nm) on Feb. 25 with the 3-m Shane reflector at the Lick Observatory. Inspection of the uncalibrated data reveals that SN 1995G is a peculiar type II supernova bearing a close resemblance to SN 1994W shortly after discovery (cf. IAUC 6046, 6057). Strong, relatively narrow (FWHM 1200 km/s), hydrogen Balmer emission lines are superposed on broader bases (FWZI 7000 km/s). Narrow, weaker Fe II emission lines are also present. The Balmer lines exhibit narrow (FWHM 300 km/s) P-Cyg absorption components with minima displaced by ∼ 700 km/s from the emission-line cores. Similar features are seen in the Fe II multiplets; they are probably produced by a dense circumstellar shell.

COMET 7P/PONS-WINNECKE CCD magnitude estimates: 1994 Dec. 31.46 UT, R = 20.1 (S. M. Lar-

son and C. W. Hergenrother, Catalina, 1.5-m reflector); 31.49, V = 20.9 (J. V. Scotti, Kitt Peak, 0.9-m Spacewatch telescope; stellar appearance); 1995 Jan. 7.49, V = 20.6 (Scotti); 28.30, R = 20.7 (Larson and Hergenrother, Kitt Peak, 2.3-m reflector).

1995 February 25 Brian G. Marsden